OpenStudy (anonymous):

medal for help - write a function rule for the area a of a triangle whose base , b is 2 cm less than 7 times the height.what is the area of the triangle when the height is 14 cn? Choices below :)

OpenStudy (dbzfan836):

hmm... 14*7-2

OpenStudy (dbzfan836):

C

OpenStudy (dbzfan836):

a=7h-2

geerky42 (geerky42):

\[A=\dfrac{1}{2}bh=\dfrac{1}{2}(\color{red}{7h-2})h\]

geerky42 (geerky42):

lol C is not right...

OpenStudy (anonymous):

oh no ?

geerky42 (geerky42):

do you understand how you plug (7h-2) into b?

OpenStudy (anonymous):

is it a?

geerky42 (geerky42):

yep it's A

OpenStudy (anonymous):

@geerky42

OpenStudy (dbzfan836):

can you explain, @geerky42

OpenStudy (triciaal):

b = 7h-2 Area, A = 1/2* b*h A = 1/2 *(7h-2)* h

geerky42 (geerky42):

b=7h-2, correct? you just plug it in formula for area of triangle A = (1/2) bh and simplify.

OpenStudy (dbzfan836):

now we are all confused, @mathmate

OpenStudy (triciaal):

distribute the h to get (7h^2 -2h) /2 answer choice A

OpenStudy (dbzfan836):

what about the ^2? that isnt supposed to be there

OpenStudy (dbzfan836):

b=7h-2 a=.5*[bh]

OpenStudy (dbzfan836):

a=bh/2

geerky42 (geerky42):

... \(\dfrac{1}{2}(7h-2)h\) simplify that @Dbzfan836 . what do you get?

OpenStudy (anonymous):

٩◔̯◔۶

OpenStudy (dbzfan836):

96/2, 48; 48*14

OpenStudy (anonymous):

head spinning

OpenStudy (triciaal):

when the height is 14 cm it means h= 14 Area = (7h^2 -2h)/2 = 7*(14)^2 - 2*14 divided by 2 = 672

OpenStudy (dbzfan836):

woah

OpenStudy (dbzfan836):

that is one trippy question

OpenStudy (dbzfan836):

A is the answer

OpenStudy (anonymous):

a is correct but the other two you helped with are also correct @Dbzfan836 ty all

OpenStudy (dbzfan836):

np

OpenStudy (dbzfan836):

(ノಠ益ಠ)ノ彡┻━┻